### v2

# DESY 25-127 arXiv:2509.16124 [hep-ph] ZU-TH 59/25 September 2025

Abstract
We present quantitative results on the single-mass heavy-flavor contributions in the region of large virtualities $Q^{2}$ up to three-loop order to the unpolarized structure function $F_{2}(x,Q^{2})$ and the polarized structure function $g_{1}(x,Q^{2})$ for the first time. These results are relevant for precision QCD analyses of the World deep-inelastic data and the data taken at future colliders, such as the Electron–Ion Collider, since the scaling violations due to massless and massive Wilson coefficients are significantly different. In order to measure the strong coupling constant $\alpha_{s}(M_{Z}^{2})$ and the twist-2 parton distribution functions consistently at highest precision, the next-to-next-to-leading order corrections have to be taken into account. Furthermore, the complete three-loop corrections will allow to reduce the present theory error of the charm mass $m_{c}$ as measured form deep-inelastic data. We provide a fast and precise public numerical code for the unpolarized and polarized massive Wilson coefficients in the asymptotic region.
1 Introduction
The heavy-flavor corrections to deep-inelastic structure functions form an essential contribution in the region of smaller values of Bjorken $x$. Their scaling violations are quite different from those of the massless contributions, which requires their detailed knowledge in QCD precision analyses of these structure functions, such as the precision measurement of the strong coupling constant $a_{s}=\alpha_{s}/(4\pi)=g_{s}^{2}/(16\pi^{2})$ [1, 2, 3, 4], the charm quark mass $m_{c}$ [5], and the parton distribution functions, cf., e.g., Refs. [6, 7, 8].
With the final results on the next-to-next-to-leading order (NNLO) single-mass corrections to the heavy-flavor Wilson coefficients in the region $Q^{2}\gg m_{Q}^{2}$, where $m_{Q}$ is the heavy-quark mass, numerical predictions on the structure functions $F_{2}$ and $g_{1}$ can be made at this order for the first time.
As has been shown in Ref. [5], the extraction of the charm quark mass from DIS data relying on approximate representations [9] leads to a large theory uncertainty of $\delta m_{c}^{\rm theory}=\tiny{\begin{array}[]{c}+0.00\\ -0.07\end{array}}\penalty 10000\ \rm GeV$ while the experimental error is $0.03\penalty 10000\ \rm GeV$. The experimental error will even decrease with new data from the Electron-Ion Collider (EIC) in the future [10, 11]. Moreover, a series of Wilson coefficients, which are numerically smaller were not considered in this analysis. The completed NNLO calculation will lower this uncertainty.
The purpose of this paper is to compile the single-mass three-loop contributions to the heavy-flavor Wilson coefficients to quantify the effect of the different contributions prior to analyses of experimental data. In the polarized case our results prepare for future precision measurements at the EIC. These quantifications were not given before. Furthermore, a public numerical library for the massive Wilson coefficients is provided.
The massive Wilson coefficients can be represented by Mellin convolutions of the massive operator matrix elements (OMEs) and pieces of the massless Wilson coefficients. The asymptotic single-mass Wilson coefficients are given by, cf. Ref. [12],
$L_{2,q}^{\rm NS}$ $=$ $a_{s}^{2}\left[A_{qq,Q}^{{\rm NS},(2)}+\hat{C}^{{\sf NS},(2)}_{2,q}\right]+a_{s}^{3}\left[A_{qq,Q}^{{\rm NS},(3)}+A_{qq,Q}^{{\rm NS},(2)}C_{2,q}^{{\rm NS},(1)}+\hat{C}^{{\rm NS},(3)}_{2,q}\right$], (1)
${\tilde{L}}_{2,q}^{\rm PS}$ $=$ $a_{s}^{3}\left[\penalty 10000\ \tilde{A}_{qq,Q}^{{\rm PS},(3)}+A_{gq,Q}^{(2)}\penalty 10000\ \penalty 10000\ \tilde{C}_{2,g}^{(1)}+\hat{\tilde{C}}^{{\rm PS},(3)}_{2,q}\right$], (2)
${\tilde{L}}_{2,g}$ $=$ $a_{s}^{2}A_{gg,Q}^{(1)}\tilde{C}_{2,g}^{(1)}+a_{s}^{3}\Bigl[\penalty 10000\ \tilde{A}_{qg,Q}^{(3)}+A_{gg,Q}^{(1)}\penalty 10000\ \penalty 10000\ \tilde{C}_{2,g}^{(2)}+A_{gg,Q}^{(2)}\penalty 10000\ \penalty 10000\ \tilde{C}_{2,g}^{(1)}+\penalty 10000\ A_{Qg}^{(1)}\penalty 10000\ \penalty 10000\ \tilde{C}_{2,q}^{{\rm PS},(2)}+\hat{\tilde{C}}^{(3)}_{2,g}\Bigr$], (3)
$H_{2,q}^{\rm PS}$ $=$ $a_{s}^{2}\left[\penalty 10000\ A_{Qq}^{{\sf PS},(2)}+\penalty 10000\ \tilde{C}_{2,q}^{{\sf PS},(2)}\right]+a_{s}^{3}\Bigl[\penalty 10000\ A_{Qq}^{{\rm PS},(3)}+\penalty 10000\ \tilde{C}_{2,q}^{{\rm PS},(3)}+A_{gq,Q}^{(2)}\penalty 10000\ \tilde{C}_{2,g}^{(1)}+A_{Qq}^{{\rm PS},(2)}\penalty 10000\ C_{2,q}^{{\sf NS},(1)}\Bigr],$ (4)
$H_{2,g}$ $=$ $a_{s}\left[A_{Qg}^{(1)}+\tilde{C}_{2,g}\right]+a_{s}^{2}\left[A_{Qg}^{(2)}+A_{Qg}^{(1)}C_{2,q}^{{\rm NS},(1)}+A_{gg,Q}^{(1)}\tilde{C}_{2,g}^{(1)}+\tilde{C}_{2,g}^{(2)}\right]+a_{s}^{3}\left[A_{Qg}^{(3)}+A_{Qg}^{(2)}C_{2,q}^{{\rm NS},(1)}\right$. (5)
$+A_{gg,Q}^{(2)}\tilde{C}_{2,g}^{(1)}+A_{Qg}^{(1)}\left(C_{2,q}^{{\rm NS},(2)}+\tilde{C}_{2,q}^{{\rm PS},(2)}\right)\left.+A_{gg,Q}^{(1)}\tilde{C}_{2,g}^{(2)}+\tilde{C}_{2,g}^{(3)}\right]$
in Mellin space, where the Mellin convolutions simplify to ordinary products. Here $A_{ij}^{(k)}$ denote the expansion coefficients of the massive OME $\langle j|O^{i}|j\rangle$ of a local twist-2 operator $O$ located at the line $i=Q,q,g$ with the external state $j=q,g$, where $Q$ denotes a massive quark, $q$ a massless quark, $g$ the gluon. The expansion coefficients of the massless Wilson coefficients $C_{i}^{l,(k)}$ depend on $N_{F}+1$ flavors and $\tilde{f}(N_{F})=f(N_{F})/N_{F}$. For the other massive Wilson coefficients see Refs. [13, 12]. The massive OMEs have been calculated to three-loop order in the unpolarized cases in Refs. [14, 19, 15, 16, 9, 17, 18] and in the polarized cases in Refs. [19, 21, 20, 17, 18]. [^1] The one– and two–loop corrections were calculated in Refs. [22, 23, 24, 25, 26, 27] and [30, 31, 13, 28, 29, 32, 33, 34].
Beyond the massive OMEs contributing to the massive Wilson coefficients, there are also others, $(\Delta)A_{gq,Q}$ and $(\Delta)A_{gg,Q}$, playing a role in the variable flavor number scheme [35, 16, 36, 38, 39, 37].
It has been shown in Ref. [13] that for the structure function $F_{2}(x,Q^{2})$ the heavy-flavor corrections in the asymptotic region $Q^{2}\gg m_{Q}^{2}$ agree with the complete result at the level of $O(1\%)$ if $Q^{2}/m_{Q}^{2}>10$ at two-loop order. This cut also safely removes the higher-twist effects, cf. Refs. [40, 41, 42]. The structure function $F_{2}(x,Q^{2})$ consists of the massless part, $F_{2}^{\rm light}(x,Q^{2})$ for the light flavors, and the massive part, $F_{2}^{\rm heavy}(x,Q^{2})$, where the Wilson coefficients have either a charm or a bottom quark contribution. [^2] Here we calculate the ‘extrinsic’ heavy-flavor contributions, but not intrinsic heavy-flavor corrections, as suggested by a Fock-space formalism in the infinite-momentum frame [43, 44].
$F_{2}(x,Q^{2})=F_{2}^{\rm light}(x,Q^{2})+F_{2}^{c}(x,Q^{2})+F_{2}^{b}(x,Q^{2}).$ (6)
In our illustrations, we work in the fixed flavor number scheme for $N_{F}=3$ light flavors with charges $e_{k}$. The NkLO contribution to $F_{2}^{\text{total}}$ combines $F_{2}^{\text{light}}$ contributions through $O(a_{s}^{k})$ and $F_{2}^{\text{heavy}}$ contributions through $O(a_{s}^{k+1})$ for $k=0,1,2$. The formal structures are the same for the structure functions $F_{2}(x,Q^{2})$ and $g_{1}(x,Q^{2})$, and are obtained by substituting the unpolarized building blocks with the polarized ones. For $g_{1}(x,Q^{2})$ the prefactor $x$ in Eqs. (7,10) has to be replaced by $1/2$.
The light-flavor part of $F_{2}$ is given by [45, 46]
$F_{2}^{\rm light}(x,Q^{2})$ $=$ $x\sum_{k=1}^{N_{F}}e_{k}^{2}\Biggl\{\frac{1}{N_{F}}\Biggl[C_{2,q}^{\rm S}\left(x,\frac{Q^{2}}{\mu^{2}}\right)\otimes\Sigma(x,\mu^{2})+C_{2,g}^{\rm S}\left(x,\frac{Q^{2}}{\mu^{2}}\right)\otimes G(x,\mu^{2})\Biggr]$ (7)
$+C_{2,q}^{\rm NS}\left(x,\frac{Q^{2}}{\mu^{2}}\right)\otimes\Delta_{k}(x,\mu^{2})\Biggr\},$
where $\Delta_{k}(x,\mu^{2}),\Sigma(x,\mu^{2})$ and $G(x,\mu^{2})$ denote the non-singlet, singlet and gluon distribution with
$\Delta_{k}(x,\mu^{2})$ $=$ $f_{k}(x,\mu^{2})+{f}_{\bar{k}}(x,\mu^{2})-\frac{1}{N_{F}}\Sigma(x,\mu^{2}),$ (8)
$\Sigma(x,\mu^{2})$ $=$ $\sum_{k=1}^{N_{F}}[f_{k}(x,\mu^{2})+{f}_{\bar{k}}(x,\mu^{2})],$ (9)
with $f_{k,(\bar{k})}$ the quark and antiquark number densities, and $\mu$ denotes the factorization scale, which we set equal to the renormalization scale. The unpolarized massless Wilson coefficients $C_{2,i}^{j}$ were calculated to three-loop order in Refs. [45, 46] and the polarized ones $\Delta C_{1,i}^{j}$ in Ref. [46]. The heavy-flavor part is given by, cf. Ref. [12],
$F_{2}^{Q}(x,Q^{2})=$ (10)
$x\Biggl\{\sum_{k=1}^{N_{F}}e_{k}^{2}\Biggl[L_{2,q}^{\rm NS}(x)\otimes(f_{k}(x,\mu^{2})+f_{\bar{k}}(x,\mu^{2}))+\frac{1}{N_{F}}\left({L}_{2,q}^{\rm PS}(x)\otimes\Sigma(x,\mu^{2})+{L}_{2,g}^{\rm S}(x)\otimes G(x,\mu^{2})\right)\Biggr]$
$+e_{Q}^{2}\Biggl[H_{2,q}^{\rm PS}(x)\otimes\Sigma(x,\mu^{2})+H_{2,g}^{\rm S}(x)\otimes G(x,\mu^{2})\Biggr]\Biggr\}.$
It contains both flavor non-singlet and singlet contributions. [^3] For the study of pure non-singlet structure functions, see Ref. [47]. There the heavy-flavor corrections are small. Here we have suppressed the dependence on $Q^{2},m_{Q}^{2}$ and $\mu^{2}$ in the heavy-flavor Wilson coefficients $H_{i}$ and $L_{i}$, and $\otimes$ denotes the Mellin convolution, [^4] For the evaluation of the convolution integrals in the numerical illustrations we use the integrator of Ref. [48].
$A(x)\otimes B(x)=\int_{0}^{1}dx_{1}\int_{0}^{1}dx_{2}\delta(x-x_{1}x_{2})A(x_{1})B(x_{2}).$ (11)
The following convolutions contribute, cf. Eqs. (256–258) of Ref. [34],
$\left(\left[\frac{f}{1-x}\right]_{+}\otimes g\right)(x)$ $=$ $\int_{x}^{1}dz\frac{f(z)}{1-z}\left[\frac{1}{z}g\left(\frac{x}{z}\right)-g(x)\right]-g(x)\int_{0}^{x}dz\frac{f(z)}{1-z},$ (12)
$(h\otimes g)(x)$ $=$ $\int_{x}^{1}\frac{dz}{z}h(z)g\left(\frac{x}{z}\right),$ (13)
$(\delta(1-x)\otimes g)(x)$ $=$ $g(x).$ (14)
Here $f(x)$ and $g(x)$ are functions with support $x\in]0,1[$.
Figure 1: The structure function $F_{2}(x,Q^{2})$ at NNLO.
In Mellin space, the massive Wilson coefficients are polynomials of the expansion coefficients of the massive OMEs $(\Delta)A_{ij}$ and expansion coefficients of the massless Wilson coefficients, see Refs. [12, 16]. The massless Wilson coefficients to three-loop order can all be expressed in terms of either harmonic sums [49, 50] in Mellin space or harmonic polylogarithms [51] in $x$-space. For the massive Wilson coefficients, this only applies to the $N_{F}$-contributions [14] and to $(\Delta)L_{q}^{\rm NS},(\Delta)L_{q}^{\rm PS}$ and $(\Delta)L_{g}^{\rm S}$. The pure-singlet Wilson coefficients $(\Delta)H_{q}^{\rm PS}$ depend on generalized harmonic sums [52, 53] in Mellin space and can be cast into harmonic polylogarithms with changed argument in $x$-space. The gluonic Wilson coefficients $(\Delta)H_{g}^{\rm S}$ receive also finite binomial sum [54] and higher transcendental function contributions in Mellin space. They contain square-root valued iterated integrals [54] and $_{2}F_{1}-$solutions [55] in $x$-space. [^5] For surveys on the different mathematical structures and the used computer algebraic methods see Refs. [56, 57]. Since we calculate the heavy-flavor corrections to inclusive structure functions, there are also contributions with massless final states, but virtual heavy-flavor corrections.
The paper is organized as follows. In Section 2 we compute the single-mass heavy-quark corrections for the unpolarized structure function $F_{2}(x,Q^{2})$ and in Section 3 those for the polarized structure function $g_{1}(x,Q^{2})$ to three-loop order. The details of the fast and precise numerical code provided, which is well suited for QCD-analyses of hard-scattering data, are described in Section 4 and the conclusions are contained in Section 5.
2 The Structure Function $F_{2}(x,Q^{2})$
In the unpolarized case, we use the parton distribution functions (PDFs) of Ref. [7] for $N_{F}=3$. [^6] From LHAPDF [58].
Figure 2: The different heavy-flavor contributions due to charm quarks to the structure function $F_{2}(x,Q^{2})$ by the Wilson coefficients $H_{g}^{\rm S},H_{q}^{\rm PS},L_{q}^{\rm NS},L_{g}^{\rm S}$ and $L_{g}^{\rm PS}$ at $Q^{2}=100\penalty 10000\ \rm GeV^{2}$ at different orders in the strong coupling constant up to $O(a_{s}),O(a_{s}^{2})$ and $O(a_{s}^{3})$.
For the values of $\alpha_{s}$ in the Wilson coefficients, we use flavor matching with the values $\alpha_{s}(25\penalty 10000\ \rm GeV^{2})=0.2020,\alpha_{s}(10^{2}\penalty 10000\ \rm GeV^{2})=0.1706$, and $\alpha_{s}(10^{3}\penalty 10000\ \rm GeV^{2})=0.1359$, corresponding to $\alpha_{s}(M_{Z}^{2})=0.1147$, consistent with the PDF fits. The massive Wilson coefficients were calculated referring to on-shell heavy-quark masses. We use the following values
$m_{c}=1.59\penalty 10000\ \rm GeV,\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ {\it m_{b}}=4.78\penalty 10000\ \rm GeV,$ (15)
cf. Refs. [5, 59].
In Figure 1, we illustrate the prediction for the total structure function $F_{2}^{\rm total}(x,Q^{2})$, including the single-mass charm and bottom quark contributions. In the region $x\sim 10^{-5}$, it rises from $Q^{2}=25\penalty 10000\ \rm GeV^{2}$ to $Q^{2}=1000\penalty 10000\ \rm GeV^{2}$ from values of $\sim 3.5$ to 10.5.
Figure 2 shows the contributions of the different heavy-flavor Wilson coefficients to $F_{2}^{c}$ at $Q^{2}=100\penalty 10000\ \rm GeV^{2}$. The different lines illustrate also the contributions from $O(a_{s})$ to $O(a_{s}^{3})$ to the structure function. The largest contribution is due to the gluonic Wilson coefficient $H_{g}^{\rm S}$ which was scaled down by a factor of 20 for better readability. The corrections are positive and grow with the order in $a_{s}$. It is followed in size by a negative contribution from the pure-singlet Wilson coefficient $H_{q}^{\rm PS}$, which starts at $O(a_{s}^{2})$. In the large-$x$ region, the non-singlet Wilson coefficient $L_{q}^{\rm NS}$ yields the dominant part.
Figure 3: The heavy-flavor contributions due to charm and bottom quarks to the structure function $F_{2}(x,Q^{2})$ at NNLO.
Around $x=0.1$ there is a turning point above which the structure function takes lower values for larger values of $Q^{2}$. However, the corrections compared to $F_{2}^{c,b}$ stay rather small. The corrections can become negative since virtual effects of heavy quarks with massless final states are contained in $F_{2}^{c}$ and $F_{2}^{b}$ and some Wilson coefficients start at $O(a_{s}^{2})$ only. Also the contributions of $L_{g}^{\rm S}$ and $L_{g}^{\rm PS}$ are of importance at the level of accuracies of $O(1\%)$. $L_{g}^{\rm S}$ is tiny at $O(a_{s}^{2})$, but receives a relative large correction at $O(a_{s}^{3})$. A part of Wilson coefficients emerging at higher order in the coupling constant $a_{s}$ can be negative as long as the unpolarized structure function remains positive.
In Figure 3 the heavy-flavor corrections to $F_{2}(x,Q^{2})$ are shown at NNLO. At $x=10^{-5}$, they grow from 1 to $\sim 4$ for $Q^{2}=25\penalty 10000\ \rm GeV^{2}$ to $Q^{2}=1000\penalty 10000\ \rm GeV^{2}$. In the large-$x$ region the corrections are negative.
The relative single-mass charm and bottom contributions to $F_{2}(x,Q^{2})$ are shown in Figure 4, as the ratio
$R=\frac{F_{2}^{c}+F_{2}^{b}}{F_{2}^{\rm light}+F_{2}^{c}+F_{2}^{b}},$ (16)
at NNLO. In the small-$x$ region of $x\sim 10^{-5}$, the fractions of the heavy-flavor corrections evolve from $Q^{2}=25\penalty 10000\ \rm GeV^{2}$ to $Q^{2}=1000\penalty 10000\ \rm GeV^{2}$ from 26 % to 41 %. At $x\sim 0.2$, the heavy-flavor corrections become zero, which is due to a combination of real and virtual corrections at NLO and NNLO. In the very-large-$x$ region, the heavy-flavor contributions at NNLO become negative and as large as $10\%$, but the structure function $F_{2}$ itself remains non-negative. It is clearly seen that the scaling violations of the massive contributions are rather different compared to those in the massless case up to high values of $Q^{2}$.
Figure 4: The ratio of the heavy-flavor contributions due to charm and bottom quarks to the structure function $F_{2}(x,Q^{2})$ at NNLO.
3 The Structure Function $g_{1}(x,Q^{2})$
In the polarized case the calculation is performed in the Larin scheme [60]. Both for the massive three-loop OMEs and the massless three-loop Wilson coefficients, the scheme transformations to the $\overline{\rm MS}$ scheme are not yet known. However, if all contributing quantities, including also the
Figure 5: The structure function $g_{1}(x,Q^{2})$ at NNLO.
Figure 6: The different heavy-flavor contributions due to charm quarks to the structure function $g_{1}(x,Q^{2})$ due to the Wilson coefficients $\Delta H_{g}^{\rm S},\Delta H_{q}^{\rm PS},\Delta L_{q}^{\rm NS},\Delta L_{g}^{\rm S}$ and $\Delta L_{g}^{\rm PS}$ at $Q^{2}=100\penalty 10000\ \rm GeV^{2}$ in the strong coupling constant up to $O(a_{s}),O(a_{s}^{2})$ and $O(a_{s}^{3})$.
parton distribution functions, are used in this scheme, $g_{1}(x,Q^{2})$, as a scheme independent quantity, can still be assembled.
For the process of deep-inelastic scattering, the Larin scheme is a consistent scheme, given the number of contributing $\gamma_{5}$ Dirac matrices. A set of Larin-scheme parton distribution functions has been generated in Ref. [61] and is used in the following illustrations. The massless and massive two-loop corrections for the structure function $g_{1}(x,Q^{2})$ were calculated in Ref. [34].
In Figure 5, we illustrate the scale evolution of the complete structure function $g_{1}^{\rm total}(x,Q^{2})$ from $Q^{2}=25\penalty 10000\ \rm GeV^{2}$ to $Q^{2}=1000\penalty 10000\ \rm GeV^{2}$ at three-loop order. With growing values of $Q^{2}$, $xg_{1}^{\rm total}(x,Q^{2})$ grows below $x\sim 0.2$ and depletes at larger values of $x$. Also below $x\sim 10^{-2}$ the order in the scales reverts. The structure functions $g_{1}^{p,d}(x,Q^{2})$ are positive within errors, cf. Ref. [62].
Figure 6 illustrates the contributions of the different heavy-flavor Wilson coefficients to $g_{1}^{c}$ for $Q^{2}=100\penalty 10000\ \rm GeV^{2}$, also with their contributions up to a given order in $a_{s}$. One clearly sees the importance of the higher-order corrections. The largest contribution is due to the gluonic Wilson coefficient $\Delta H_{g}^{\rm S}$. As has been discussed in Refs. [32, 34], the first moment of $\Delta H_{g}^{\rm S}(x,Q^{2})$ vanishes,
$\int_{0}^{1}dx\Delta H_{g}^{\rm S}(x,Q^{2})=0.$ (17)
Figure 7: The heavy-flavor contributions due to charm and bottom quarks to the structure function $g_{1}(x,Q^{2})$ at NNLO.
This is now also found at three-loop order. Mainly due to this, the oscillating structure in Figure 6 is implied. Next in size are the flavor non-singlet and pure-singlet Wilson coefficients, $\Delta L_{q}^{\rm NS}$ and $\Delta H_{q}^{\rm PS}$. In the small-$x$ region, the former is positive and the latter negative. $\Delta L_{q}^{\rm NS}$ turns negative above $x\sim 0.2$ and dominates the large-$x$ region. The contributions due to $\Delta L_{g}^{\rm S}$ and $\Delta L_{g}^{\rm PS}$ are smaller but not negligible.
Figure 7 shows the NNLO heavy-flavor contributions to $xg_{1}(x,Q^{2})$. They grow in size with $Q^{2}$. Below $x\sim 0.01$ they are negative and turn positive until at $x\sim 0.3$ they turn negative again.
In Figure 8, we show the ratio Eq. (16) of the charm and bottom contributions to the whole structure function $g_{1}(x,Q^{2})$ at NNLO for $Q^{2}=25,100$ and $1000\penalty 10000\ \rm GeV^{2}$. In the small-$x$ region, the corrections are negative. The largest negative values are about $10\penalty 10000\ \%,-25\penalty 10000\ \%$ and $95\penalty 10000\ \%$ from $Q^{2}=25,100$ to $1000\penalty 10000\ \rm GeV^{2}$. These large corrections occur in a region in which $xg_{1}(x,Q^{2})$ itself is rather small, see Figure 5. Around $x=10^{-2}$, the ratio turns positive. At $x=0.3$ it turns negative again, as in the unpolarized case. The complete structure function $g_{1}(x,Q^{2})$ remains positive. The strong negative correction in the small-$x$ region is mainly induced by the sum rule of the vanishing first moment of $\Delta H_{g}^{\rm S}$.
Figure 8: The ratio of the heavy-flavor contributions due to charm and bottom quarks to the structure function $g_{1}(x,Q^{2})$ at NNLO.
4 The numerical code
We provide the Fortran-library WILS3HQ of the massive Wilson coefficients to three-loop order as a fast and precise numerical implementation. It is well suited for QCD analyses of deep-inelastic structure functions in the unpolarized and polarized case. The Wilson coefficients are represented as polynomials in the parameters $a_{s},N_{F}$, $\ln(Q^{2}/\mu^{2}),\ln(m^{2}/\mu^{2})$ and flav, see below. Also the $x$-dependence is described by elementary functions and polynomial interpolations. The order in the coupling constant is selected by the parameter IO = 1,2,3. Both for the structure functions $F_{2}$ and $g_{1}$ five different Wilson contribute. In the non-singlet cases $L_{q,2(1)}^{\rm NS}$ there are two distribution valued contributions in addition to the regular part. Likewise, for the Wilson coefficients $H_{2(1),q}^{\rm PS}$ and $L_{2(1),q}^{\rm PS}$, $\delta(1-x)$-terms contribute. The regular terms are represented by suitably deep analytic small- and large-$x$ expansions and grid interpolations for the remainder parts for $x\in[0,1]$, optimized for $x\geq x_{\rm min}=10^{-6}$. The grid interpolations, performed by third order splines [63], are based on 501 points. [^7] We thank S. Kumano and M. Miyama of the AAC-collaboration for allowing us to use their interpolation routines. The provided spline representation can be addressed directly by setting IDAT = 1. The ++-and $\delta$-distribution parts are given in analytic form. The representations need only elementary functions and we used code optimization [64]. The Mellin convolutions with the parton distributions have to be performed in the respective implementations by the users. The data-set initialization is done by the routines INIDAT1 ... 10, where the order of the Wilson coefficients is H2gS, H2qPS, L2qNS, L2gS, L2qPS; HG1gS, HG1qPS, LG1qNS, LG1gS, LG1qPS.
For the grid-files we use the following name convention, illustrated by the example
${\tt L2qPS30011.dat}.$ (18)
The name starts with the name of the Wilson coefficient. The number sequence denotes the order in $a_{s}$, the power of $\ln(m^{2}/\mu^{2})$, the power of $\ln(Q^{2}/\mu^{2})$, the presence of the factor $N_{F}$, the presence of the factor $\tt flav$. The set in Eq. (18) corresponds to the grid for the massive Wilson coefficient $L_{2q}^{\rm PS(3)}$ of the structure function $F_{2}$, proportional to the factor $a_{s}^{3}N_{F}{\tt flav}$.
charm bottom
$L_{q}^{{\rm NS}}$ 22 1-1
$L_{q}^{{\rm PS}}$ $\frac{2}{5}$ $\frac{2}{7}$
$L_{g}^{{\rm S}}$ $\frac{2}{5}$ $\frac{1}{7}$
$H_{q}^{{\rm PS}}$ $\frac{2}{5}$ $\frac{2}{7}$
$H_{g}^{{\rm S}}$ $\frac{2}{5}$ $\frac{1}{7}$
Table 1: Numerical values of the factor flav for the perturbative generation of a heavy charm or bottom quark.
For the charm quark we find the following formulae for the flavor factors [^8] These qunatities are related to the numbers ${\sf w}_{2}$ and ${\sf w}_{3}$ in Refs. [45, 46].
[…]
with $e_{c}$ the electric charge of the charm quark and we already inserted the charges of the three light quark flavors. The corresponding numbers in the case of bottom are obtained by replacing $e_{c}$ with $e_{b}$. The flavor factors for $L_{q}^{\rm PS}$ and $H_{q}^{\rm PS}$ as well as $L_{g}^{\rm S}$ and $H_{g}^{\rm S}$ are the same, since all flavor factors vanish when considering only the three light quark charges. These flavor factors result from the pieces of the three-loop massless Wilson coefficients in the asymptotic massive Wilson coefficients.
The regular parts of the unpolarized Wilson coefficients are denoted by REG...f, as, e.g., REGL2qNS.f in the flavor non-singlet Wilson coefficient for the structure function $F_{2}$ and, correspondingly, with REGLG1qNS.f for the structure function $g_{1}$. The routines of the ++-distribution parts are named by PLU...f, and the coefficients of the $\delta(1-x)$ terms by DEL...f. The argument list is
${\tt(IO,X,as,LM,LQ,NF,flav)},$ (24)
where ${\tt as}=a_{s}(\mu^{2}),{\tt LM}=\ln(m^{2}/\mu^{2}),{\tt LQ}=\ln(Q^{2}/\mu^{2}),{\tt NF}=N_{F},{\tt X}=x$, the Bjorken variable. In the case of the $\delta(1-x)$ terms X is not part of the argument list. For the regular parts there are the auxiliary functions SX....f, LX....f, and GR....f, for the small- and large-$x$ expansions, and the grid part.
The different functions of the library can be compiled with gfortran. We provide ten programs MAIN1.f to MAIN10.f to test the set-up of the respective Wilson coefficients by comparing with numerical results obtained by our initial (semi)-analytic calculation, for some specific choice of parameters.
We add the numerical integration routine DAIND of Ref. [65] to perform the convolution integrals with the parton densities for the convenience of the user. The license conditions to use the library WILS3HQ are to quote the papers Refs. [12, 14, 19, 15, 16, 20, 21, 17, 18] and the present paper, in which the single-mass three-loop corrections were calculated.
5 Conclusions
In this paper, we present for the first time numerical results for the single-mass three-loop heavy-flavor corrections to the unpolarized structure function $F_{2}(x,Q^{2})$ and the polarized structure function $g_{1}(x,Q^{2})$ at twist-2. The asymptotic analytic results on the heavy-flavor Wilson coefficients are correlated to the kinematic region in which the higher-twist contributions can be neglected. Our numerical results illustrate the impact of the different Wilson coefficients and show the changes in the structure functions going from the $O(a_{s})$ to the $O(a_{s}^{3})$ corrections. In the unpolarized case, the corrections are large in the small-$x$ region, which is also observed in the polarized case. Here, however, the large corrections emerge in a region where the structure function is small, unlike in the former case. The present results are an important step to allow for consistent QCD analyses of the deep-inelastic World data at the level of NNLO, given the partial results used before for combined singlet/non-singlet data analyses. Our new results will allow to improve the accuracy of the strong coupling constant $a_{s}(M_{Z}^{2})$, the PDFs and the charm mass, extracted from the deep-inelastic World data.
A fast and numerically precise public Fortran code for the calculation of the asymptotic single-mass heavy-flavor Wilson coefficients, WILS3HQ, for the use in experimental data analyses, is provided as an ancillary file to this paper.
Acknowledgments. We would like to thank I. Bierenbaum, A. Goedicke (né Hasselhuhn), S. Klein, C. Raab, M. Round, M. Saragnese and F. Wißbrock for collaboration in earlier projects and thank P. Marquard for discussions. This work was supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme grant agreement 101019620 (ERC Advanced Grant TOPUP), the UZH Postdoc Grant, grant no. [FK-24-115] and by the Austrian Science Fund (FWF) Grant-DOI 10.55776/P20347.